OpenStudy (anonymous):

One of the factors of p5 – 144p3 is: p3 p + 12 p – 12

jimthompson5910 (jim_thompson5910):

Hint: p^5 = p*p*p*p*p and p^3 = p*p*p

jimthompson5910 (jim_thompson5910):

What is the GCF here?

OpenStudy (anonymous):

all three are correct.

jimthompson5910 (jim_thompson5910):

yes, because p^5 - 144p^3 p^3(p^2 - 144) p^3(p+12)(p-12)